OpenStudy (anonymous):

im missing a bit of algeba here. . . integrate 3 / (2-sqrt x) if i use u = sqrt x then du is 1/(2 sqrt x) when wolfram does the sub it writes the integral as 6 integral u / (2-u) i understand everything except the u in the numerator. how is that u in the numerator getting there

OpenStudy (amistre64):

\[\int \frac{3}{2-\sqrt(x)}dx\]

OpenStudy (amistre64):

this is it right?

OpenStudy (amistre64):

if we claim that u = sqrt(x), then du = 1/2sqrt(x) dx; and since sqrt(x) IS u, we can rewrite this as, 1/2u ... if that helps

OpenStudy (amistre64):

so lets apply this; du = 1/2u dx 2u du = dx

OpenStudy (amistre64):

now replace what gets replaced

OpenStudy (anonymous):

the u is in the denominator because we let u=sqrtx which is being substituted in the denominator

OpenStudy (amistre64):

int 3 / 2-sqrt(x) dx int 3 * 2u / 2-u du

OpenStudy (amistre64):

\[\int 6\frac{u}{2-u}du\]

OpenStudy (amistre64):

all were missing now is a negtive and we are good to go

OpenStudy (anonymous):

3 / (2-sqrt x) dx, u=x^(1/2) so du=(1/2)x^(-1/2)dx so dx=2x^(1/2)du and if you substitute this in here 3 / (2-sqrt x) dx, you get (3/(2-u))(2x^(1/2))du, since u=x^(1/2) then (3/(2-u))(2x^(1/2))du becomes (3/(2-u))(2u)du which is (6u/2-u)du..this should answer your question..please reply back if you understand this thanks

OpenStudy (amistre64):

that could have been formated better ...

OpenStudy (amistre64):

from above .... 3 / (2-sqrt x) dx u=x^(1/2) ; du=(1/2)x^(-1/2)dx dx=2x^(1/2)du if you substitute this in here: 3 / (2-sqrt x) dx, you get: (3/(2-u))(2x^(1/2))du, since u=x^(1/2) then (3/(2-u))(2x^(1/2))du becomes (3/(2-u))(2u)du which is (6u/2-u)du.. this should answer your question..please reply back if you understand this thanks....

OpenStudy (anonymous):

that helps very much. Thank you both. Sorry I dropped out yesterday. I was trying the site for a tablet and it was not working so well. Thank you again.